# Cosmological singularity, conformal anomaly and symmetric polynomials

### v1

@@ 4 Integrating conformal anomaly
where ${\cal M}(a,b,c)$ is defined in (3.8). Resolving equation (4.17) we represent the components of stress tensor as follows
@@
Due to the permutational symmetry of the metric (2.1), it is evident that $G(a,b,c)$ is a symmetric function of the parameters $a$, $b$ and $c$. Additionally, the function $F_{1}(a,b,c)$ is symmetric with respect to the permutation of $b$ and $c$. The function $F_{2}(a,b,c)$ is symmetric with respect to $a$ and $c$, while $F_{3}(a,b,c)$ is symmetric with respect to $a$ and $b$.
@@
In the semiclassical Einstein equations modified by the contribution from a quantum conformal field theory (CFT), the left-hand side (due to the Ricci tensor) grows as $1/t^{2}$ for small $t$, while the right-hand side, arising from the conformal anomaly, grows as $1/t^{4}$. The latter therefore dominates and must vanish in any consistent solution. To leading order, the Einstein equations thus reduce to a set of algebraic equations involving two functions of the parameters $(a,b,c)$:
@@ 5 Kasner CFT
Both sides of equation (4.22) represent a symmetric polynomial of degree 5 of three variables $a,b,c$. Any such polynomial can be decomposed over the symmetric basis $e_{1},e_{2},e_{3}$, there are 16 terms in such a polynomial. Thus, equation (4.22) is equivalent to a system of 16 linear equations. On the other hand, $F(a;b,c)$ is a polynomial of degree 4 symmetric in $b$ and $c$ and there may appear 22 terms (and respective coefficients to be determined) in such a polynomial, see equation (2.1) of appendix. The naive counting thus suggests that there are 22-16=6 that remain undetermined. The direct analysis (see appendix A), however, shows that there remain 7 undetermined coefficients, so that one of the 22 equations follows from the others. In appendix A we solve equation (4.22) and find the respective coefficients as functions of these 7 coefficients that remain undetermined (notice, however, that the choice of them is rather ambiguous).
@@
This relation implies that, provided that (ab)(bc)(ca)(a-b)(b-c)(c-a) is non-vanishing, then equations $F(a;b,c)=F(b;a,c)=F(c;b,a)=0$ are equivalent to vanishing of three symmetric polynomials:
@@
The relations (5.4), (5.6), (5.7), (5.8) completely solve the problem of integrating the conformal anomaly in a Kasner type metric (2.1) and define the components of the CFT stress tensor according to equation (4.19).
@@ 6 Solving the constraints
Thus, in the previous section we completely integrated the conformal anomaly for a Kasner type metric (2.1), i.e. we have found the respective stress energy tensor that is covariantly conserved and its trace produces correctly the anomaly. Now, we want to solve the respective complete gravitational equations considering the anomalous stress tensor as a source for the gravitational field. These equations are tensorial. They include vanishing of the anomaly, in the leading oder in $t$, as one of the constraints. As we have already explained in the previous sections, the equations of semiclassical gravity, to leading order, reduce to vanishing of three symmetric functions, $K(a,b,c)=0$, $S(a,b,c)=0$ and $H(a,b,c)=0$, where the functions are given be equations (5.6), (5.7) and (5.8). They are defined up to 7 undetermined parameters, $k_{1}$, $k_{3}$, $m_{1}$, $m_{2}$, $m_{3}$, $m_{4}$, $m_{5}$. If all these coefficients are non-vanishing then equation $K(a,b,c)=0$ is solved by expressing $e_{2}$ as a quadratic function of $e_{1}$ and then it should be substituted into equation $S(a,b,c)=0$ that is solved by expressing $e_{3}$ as a cubic function of $e_{1}$. All this should be then substituted into equation $H(a,b,c)=0$ that becomes a quartic equation for $e_{1}$. This equation may have at most 4 real solutions thus expressing $e_{1}$ as function of the 7 coefficients. Thus, in this case the solution of all three equations in the space of parameters $(a,b,c)$ is represented by four isolated points the position of which is determined by the values of the seven coefficients.
@@
As expected, the system of equations $S(a,b,c)=0$ and $H(a,b,c)=0$ is equivalent to Master equation ${\cal M}(a,b,c)=0$ (3.11). Indeed, equation $S(a,b,c)=0$ can be used to express $m_{2}$ in terms of $e_{1}$ and $e_{2}$ that can be substituted in equation (6.3), one then obtains
@@
In the case when $\eta=1$ this equation reduces to equation $(e_{1}-1)^{2}y(a,b,c)=0$. An obvious solution is $e_{1}=1$ (and $e_{2}=0$). This was our choice of the coefficients in equations (5.7) and (5.8) to have this classical Kasner solution of the semiclassical equations when $\eta=1$. For $e_{1}$ different from $1$ the equations reduce to $y(a,b,c)=0$. This is cubic equation. Thus, it may have at most 3 real solutions for a given $m_{2}$. For each value of $e_{1}$ the corresponding value of $e_{2}$ then is given by equation (6.5). Value of $e_{3}$ remains arbitrary in this case. This is in full analogy with what we had in the classical case.
@@
Singular but geodesically complete metrics as solutions. It is instructive to examine more closely the limit of large negative values of $e_{1}$. This regime is reached for small negative values of $m_{2}$, since $e_{1}=\frac{12}{m_{2}}(1-\frac{1}{4}m_{2}-\frac{1}{12}m_{2}^{2}+O(m_{2}^{3}))$ as can be seen from equation $y(a,b,c)=0$. This is blue curve labelled B in Figure 8. From equation (6.5) we find that $e_{2}=\frac{e_{1}^{2}}{4}$ to the leading order, and the consistency bound is clearly satisfied. As $e_{3}$ may take any permitted value we choose $e_{3}=se_{1}^{3}$, where $s>0$ is a parameter.Then values of $a,b,c$ are three roots of cubic equation $x^{3}-e_{1}x^{2}+e_{2}x-e_{3}=0$. This equation has three real roots $x_{1}<x_{2}<x_{3}$ provided $s<1/54$. All roots are negative: $x_{1}<e_{1}/2$, $x_{2}<e_{1}/6$, $x_{3}<0$. For the third root we have an estimate that $x_{3}\simeq 0,045e_{1}$ provided $s=1/108$. Hence, for large negative $e_{1}$ (small negative $m_{2}$) all three parameters $a,b,c$ that enter the Kasner type metric (2.1) take negative values that satisfy the condition (2.9). The corresponding metric therefore possesses a curvature singularity at $t=0$; however, this singularity lies at an infinite value of the affine parameter for any null geodesic traced backward in time toward it. In other words, any light emitted from the singularity can never reach any observer at a later time. This result demonstrates that within the space of solutions to the semiclassical gravitational equations, there exists a domain of parameters in which the spacetime exhibits a curvature singularity while remaining geodesically complete.
@@
The conditions $e_{2}>0$ and $e_{2}\leq\frac{e_{1}^{3}}{3}$ are satisfied for $\lambda>3/2$. Then $e_{3}$ scales as $e_{1}^{3}$ for large negative values of $e_{1}$, the coefficient of proportionality is positive if $\lambda>3$ (for $\eta>1$) or $3/2<\lambda<3$ (for $\eta<1$). The bound (2.14) imposes extra constraint on the pre-factor in $e_{3}$ (6.11). Under these conditions all three parameters $a,b,c$ take negative values that scale linearly with $e_{1}$ and hence satisfy the condition (2.9). Whether all three parameters $a,b,c$ can simultaneously be negative and less than minus one for other values of $m_{2}$ requires further investigation.
7 Integration of anomaly in isotropic case: regular spacetime
In the previous section we mostly focused on the case when all three parameters $a,b,c$ take different values. Now we want to discuss the isotropic case when all three parameters have the same value, $a=b=c$. In section 3.3 we have already considered this case when discussed the condition of vanishing the anomaly, ${\cal M}(a,a,a)=18\eta a^{3}(a-1)$. There we concluded that this condition selects two possible value $a=0$ ad $a=1$. The first is just regular spacetime that looks as Minkowski and the second has a curvature singularity.
@@ 7 Integration of anomaly in isotropic case: regular spacetime
Let us return to the equations (4.21) and (4.22) for the functions $G(a;b,c)$ (related to component $T^{0}_{0}$ of the anomalous stress tensor) and $F(a;b,c)$ (related to the component $T^{a}_{a}$). As we discussed in sections 4, 5 and 6 solving these equations involve much freedom. In the case when $a,b$ and $c$ take same value this freedom disappears and the result is unique. Indeed, the equations (4.21) and (4.22) become equations
@@
where we use notations $G(a)=G(a;a,a)$ and $F(a)=F(a;a,a)$. This has unique solution,
@@
We stress that this solution is unambiguous: no free constants are involved. The gravitational equations then reduce to $F(a)=0$ and $G(a)=0$. We find that only value $a=0$ remains as a valid solution while $a=1$ appears to be spurious. Thus, only the isotropic spacetime, which is free of singularities, satisfies the complete semiclassical equations. This is, of course, a well-known feature of the Starobinsky model [4]. A detailed analysis of the isotropic case is presented in Appendix B, where we show that the absence of the singularity arises from a particular choice of quantum state. The still remaining freedom is in the solution of the homogeneous differential equation that we systematically neglected in the paper. In the isotropic case adding this solution is equivalent to adding an isotropic radiation to the quantum state. This generically leads to appearance of the singularity. On the other hand, as was explained in [4], the absence of such radiation in the quantum state leads to a completely regular isotropic solution of gravitational equations.
@@ 8 Conclusions
In the present paper, we address the problem of cosmological singularities in semiclassical gravity modified by a quantum conformal field theory (CFT). For Kasner-type metrics, the leading-order behavior near the singularity reduces the problem to a set of algebraic constraints imposed on the Kasner parameters. We determine the complete asymptotic structure of the anomalous stress-energy tensor of the quantum CFT, an uncommon situation, as such analyses can be carried out explicitly only in a few very special cases. Integrating the conformal anomaly for Kasner-type metrics allows us to obtain the full set of constraints on the parameters, which can be solved conveniently when expressed in terms of symmetric polynomials.
@@ Acknowledgements
Appendix A Appendix: integrating anomaly in Kasner type metric
@@
This representation has to be substituted into equation (4.22). This equation is symmetric in $a,b$ and $c$ and thus can be decomposed over basis $e_{1}$, $e_{2}$ and $e_{3}$ defined earlier in (2.10). This equation is a polynomial of degree 5 and thus has 16 independent terms. That means that equation (4.22) reduces to 16 different equations on coefficients $f_{nl}$. In (A) there are 22 such coefficients. A naive counting would suggest that there remain 22-16=6 undetermined coefficients. The direct calculation however shows that there are 7 such coefficients. It appears that one of the equations follows from the others. There is ambiguity which coefficients are to be taken as these 7 coefficients that define all other coefficients. Here is one possible choice to solve the system of linear equations that come from equation (4.22):
@@
From the 4th order polynomial $F(a;b,c)$ one may define a 3rd order polynomial $E(c;a,b)=E(c;b,a)$ by relation
@@
and then yet a 2nd order polynomial $K(b;a,c)$ defined as follows
@@
It is an important fact that for $F(a;b,c)$ defined by (A.1)-(A) the polynomial $K(b;a,c)$ is symmetric in $(a,b,c)$ so that $K(b;a,c)=K(a,b,c)$. Additionally, there is a constraint on coefficients in decomposition over a symmetric basis, the exact form then reads
@@
Appendix B Appendix: isotropic cosmology
@@
where the second term is a solution to the homogeneous equation, when the r.h.s. in (B.6) vanishes, $C_{0}$ is an integration constant. The freedom to choose $C_{0}$ is due to the choice of the quantum state. Non-vanishing $C_{0}$ can be interpreted as due to presence of isotropic radiation. One finds that,

### v2

@@ 4 General form of anomalous stress tensor
where ${\cal M}(a,b,c)$ is defined in (3.8). Upon solving equation (4.17) we represent the components of stress tensor as follows
@@
Due to the permutation symmetry of the metric (2.1), it is evident that $G(a,b,c)$ is a symmetric function of the parameters $a$, $b$ and $c$. Additionally, the function $F_{1}(a,b,c)$ is symmetric with respect to the permutation of $b$ and $c$. The function $F_{2}(a,b,c)$ is symmetric with respect to $a$ and $c$, while $F_{3}(a,b,c)$ is symmetric with respect to $a$ and $b$.
@@
In the semiclassical Einstein equations (3.2) modified by the contribution from a quantum conformal field theory (CFT), the left-hand side (due to the Ricci tensor) grows as $1/t^{2}$ for small $t$, while the right-hand side, arising from the conformal anomaly, grows as $1/t^{4}$. The latter therefore dominates and must vanish in any consistent solution. To leading order, the Einstein equations thus reduce to a set of algebraic equations involving two functions of the parameters $(a,b,c)$:
@@ 5 Kasner CFT
Both sides of equation (4.22) represent a symmetric polynomial of degree 5 of three variables $a,b,c$. Any such polynomial can be decomposed over the symmetric basis $e_{1},e_{2},e_{3}$, there are 16 terms in such a polynomial. Thus, equation (4.22) is equivalent to a system of 16 linear equations. On the other hand, $F(a;b,c)$ is a polynomial of degree 4 symmetric in $b$ and $c$ and there may appear 22 terms (and respective coefficients to be determined) in such a polynomial, see equation (2.1) of appendix. The naive counting, thus, suggests that there are 22-16=6 coefficients that remain undetermined. The direct analysis (see appendix A), however, shows that there remain 7 undetermined coefficients, so that one of the 22 equations follows from the others. In appendix A we solve equation (4.22) and find the respective coefficients as functions of these 7 coefficients that remain undetermined (notice, however, that the choice of these coefficients is rather ambiguous).
@@
This relation implies that, provided that (ab)(bc)(ca)(a-b)(b-c)(c-a) is non-vanishing, the equations $F(a;b,c)=F(b;a,c)=F(c;b,a)=0$ are equivalent to the vanishing of three symmetric polynomials:
@@
The relations (5.4), (5.6), (5.7), (5.8) completely solve the problem of finding the stress tensor with the conformal anomaly in a Kasner type metric (2.1) and define the components of the CFT stress tensor according to equation (4.19).
@@ 6 Solving the constraints
Thus, in the previous section we completely integrated the conformal anomaly for a Kasner type metric (2.1), i.e. we have found the corresponding stress energy tensor that is covariantly conserved and its trace reproduces correctly the anomaly. Now, we want to solve the complete gravitational equations considering the anomalous stress tensor as a source for the gravitational field. These equations are tensorial. They include the vanishing of the anomaly, to leading oder in $t$, as one of the constraints. As we have already explained in the previous sections, the equations of semiclassical gravity, to leading order, reduce to the vanishing of three symmetric functions, $K(a,b,c)=0$, $S(a,b,c)=0$ and $H(a,b,c)=0$, where the functions are given be equations (5.6), (5.7) and (5.8). They are defined up to 7 undetermined parameters, $k_{1}$, $k_{3}$, $m_{1}$, $m_{2}$, $m_{3}$, $m_{4}$, $m_{5}$. If all these coefficients are non-vanishing then equation $K(a,b,c)=0$ is solved by expressing $e_{2}$ as a quadratic function of $e_{1}$ which then should be substituted into equation $S(a,b,c)=0$ that is solved by expressing $e_{3}$ as a cubic function of $e_{1}$. All this should be then substituted into equation $H(a,b,c)=0$ that becomes a quartic equation for $e_{1}$. This equation may have at most 4 real solutions thus expressing $e_{1}$ as function of the 7 coefficients. Thus, in this case the solution of all three equations in the space of parameters $(a,b,c)$ is represented by four isolated points the position of which is determined by the values of the seven coefficients.
@@
As expected, the system of equations $S(a,b,c)=0$ and $H(a,b,c)=0$ is equivalent to the Master equation ${\cal M}(a,b,c)=0$ (3.11). Indeed, equation $S(a,b,c)=0$ can be used to express $m_{2}$ in terms of $e_{1}$ and $e_{2}$ that can be substituted in equation (6.3), one then obtains
@@
In the case when $\eta=1$ this equation reduces to the equation $(e_{1}-1)^{2}y(a,b,c)=0$. An obvious solution is $e_{1}=1$ (and $e_{2}=0$). This was our choice of the coefficients in equations (5.7) and (5.8) to have this classical Kasner solution of the semiclassical equations when $\eta=1$. For $e_{1}$ different from $1$ the equations reduce to $y(a,b,c)=0$. This is cubic equation. Thus, it may have at most 3 real solutions for a given $m_{2}$. For each value of $e_{1}$ the corresponding value of $e_{2}$ then is given by equation (6.5). Value of $e_{3}$ remains arbitrary in this case. This is in full analogy with what we had in the classical case.
@@
Singular but geodesically complete metrics as solutions. It is instructive to examine more closely the limit of large negative values of $e_{1}$. This regime is reached for small negative values of $m_{2}$, since $e_{1}=\frac{12}{m_{2}}(1-\frac{1}{4}m_{2}-\frac{1}{12}m_{2}^{2}+O(m_{2}^{3}))$ as can be seen from equation $y(a,b,c)=0$. This is the blue curve labelled B in Figure 8. From equation (6.5) we find that $e_{2}=\frac{e_{1}^{2}}{4}$ to the leading order, and the consistency bound is clearly satisfied. As $e_{3}$ may take any permitted value we choose $e_{3}=se_{1}^{3}$, where $s>0$ is a parameter.Then the values of $a,b,c$ are the three roots of the cubic equation $x^{3}-e_{1}x^{2}+e_{2}x-e_{3}=0$. This equation has three real roots $x_{1}<x_{2}<x_{3}$ provided $s<1/54$. All roots are negative: $x_{1}<e_{1}/2$, $x_{2}<e_{1}/6$, $x_{3}<0$. For the third root we have an estimate that $x_{3}\simeq 0,045e_{1}$ provided $s=1/108$. Hence, for large negative $e_{1}$ (small negative $m_{2}$) all three parameters $a,b,c$ that enter the Kasner type metric (2.1) take negative values that satisfy the condition (2.9). The corresponding metric therefore possesses a curvature singularity at $t=0$; however, this singularity lies at an infinite value of the affine parameter for any null geodesic traced backward in time toward it. In other words, any light emitted from the singularity can never reach any observer at a later time. This result demonstrates that within the space of solutions to the semiclassical gravitational equations, there exists a domain of parameters in which the spacetime exhibits a curvature singularity while remaining geodesically complete.
@@
The conditions $e_{2}>0$ and $e_{2}\leq\frac{e_{1}^{3}}{3}$ are satisfied for $\lambda>3/2$. Then $e_{3}$ scales as $e_{1}^{3}$ for large negative values of $e_{1}$, the coefficient of proportionality is positive if $\lambda>3$ (for $\eta>1$) or $3/2<\lambda<3$ (for $\eta<1$). The bound (2.14) imposes an extra constraint on the pre-factor in $e_{3}$ (6.11). Under these conditions all three parameters $a,b,c$ take negative values that scale linearly with $e_{1}$ and hence satisfy the condition (2.9). Whether all three parameters $a,b,c$ can simultaneously be negative and less than 1-1 for other values of $m_{2}$ requires further investigation.
7 Semiclassical solution in isotropic case: regular spacetime
In the previous section we mostly focused on the case when all three parameters $a,b,c$ take different values. Now we want to discuss the isotropic case when all three parameters have the same value, $a=b=c$. In section 3.3 we have already considered this case when we discussed the condition of vanishing of the anomaly, ${\cal M}(a,a,a)=18\eta a^{3}(a-1)$. There we concluded that this condition selects two possible values $a=0$ ad $a=1$. The first is just regular spacetime that looks like Minkowski and the second has a curvature singularity.
@@ 7 Semiclassical solution in isotropic case: regular spacetime
Let us return to the equations (4.21) and (4.22) for the functions $G(a;b,c)$ (related to component $T^{0}_{0}$ of the anomalous stress tensor) and $F(a;b,c)$ (related to the component $T^{a}_{a}$). As we discussed in sections 4, 5 and 6 solving these equations involves considerable freedom. In the case when $a,b$ and $c$ take the same value this freedom disappears and the result is unique. Indeed, the equations (4.21) and (4.22) become equations
@@
where we use notations $G(a)=G(a;a,a)$ and $F(a)=F(a;a,a)$. This has the unique solution,
@@
We stress that this solution is unambiguous: no free constants are involved. The gravitational equations then reduce to $F(a)=0$ and $G(a)=0$. We find that only the value $a=0$ remains as a valid solution while $a=1$ appears to be spurious. Thus, only the isotropic spacetime, which is free of singularities, satisfies the complete semiclassical equations. This is, of course, a well-known feature of the Starobinsky model [4]. A detailed analysis of the isotropic case is presented in Appendix B, where we show that the absence of the singularity arises from a particular choice of quantum state. The still remaining freedom is in the solution of the homogeneous differential equation that we systematically neglected in the paper. In the isotropic case adding this solution is equivalent to adding an isotropic radiation to the quantum state. This generically leads to appearance of the singularity. On the other hand, as was explained in [4], the absence of such radiation in the quantum state leads to a completely regular isotropic solution of gravitational equations.
@@ 8 Conclusions
In the present paper, we address the problem of cosmological singularities in semiclassical gravity modified by a quantum conformal field theory (CFT). For Kasner-type metrics, the leading-order behavior in the vicinity of the singularity reduces the problem to a set of algebraic constraints imposed on the Kasner parameters. We determine the complete asymptotic structure of the anomalous stress-energy tensor of the quantum CFT, an uncommon situation, as such analyses can be carried out explicitly only in a few very special cases. Integrating the conformal anomaly for Kasner-type metrics allows us to obtain the full set of constraints on the parameters, which can be solved conveniently when expressed in terms of symmetric polynomials.
@@ Acknowledgements
Appendix A Function $F(a;b,c)$ for a Kasner type metric
@@
This representation has to be substituted into equation (4.22). This equation is symmetric in $a,b$ and $c$ and thus can be decomposed over the basis $e_{1}$, $e_{2}$ and $e_{3}$ defined earlier in (2.10). This equation is a polynomial of degree 5 and thus has 16 independent terms. That means that equation (4.22) reduces to 16 different equations for the coefficients $f_{nl}$. In (A) there are 22 such coefficients. A naive counting would suggest that there remain 22-16=6 undetermined coefficients. The direct calculation however shows that there are 7 such coefficients. It appears that one of the equations follows from the others. There is an ambiguity which coefficients are to be taken as these 7 coefficients that define all other coefficients. Here is one possible choice to solve the system of linear equations that come from equation (4.22):
@@
From the 4th order polynomial $F(a;b,c)$ one may define a 3rd order polynomial $E(c;a,b)=E(c;b,a)$ by the relation
@@
and then also a 2nd order polynomial $K(b;a,c)$ defined as follows
@@
It is an important fact that for $F(a;b,c)$ defined by (A.1)-(A) the polynomial $K(b;a,c)$ is symmetric in $(a,b,c)$ so that $K(b;a,c)=K(a,b,c)$. Additionally, there is a constraint on the coefficients in the decomposition over a symmetric basis, the exact form then reads
@@
Appendix B Isotropic cosmology
@@
where the second term is a solution to the homogeneous equation, when the r.h.s. in (B.6) vanishes, $C_{0}$ is an integration constant. The freedom to choose $C_{0}$ is due to the choice of the quantum state. Non-vanishing $C_{0}$ can be interpreted as due to the presence of isotropic radiation. One finds that,